A cart of mass m1=0.24kg moves on a linear track without friction with an initial velocity of 0.17 m/s. It collides with another cart of mass m2=0.68 kg that is initially at rest. The first cart carries a force probe that registers the magnitude of the force exerted by one cart on the other during the collision.

Find the velocity of each cart after the collision.


1
Expert's answer
2021-01-07T10:55:34-0500

Impulse:


"I=0.5(10)(14-3)(10^{-3})=0.055\frac{kgm}{s}"

"p_{1f}=p_{1i}-J\\m_1v_{1}=m_1u_1-J\\=0.24(0.17)-0.055=-0.0142\frac{kgm}{s}"

"v_1=-0.059\frac{m}{s}"

"p_{2f}=J\\m_2v_{2}=J\\0.68v_{2f}=0.055\\v_{2f}=0.082\frac{m}{s}"